# The Picard group of the Baily-Borel compactification of the moduli space of polarized K3 surfaces

### v1

@@ 6.2 Heegner divisors on the boundary
(cf. [9, Theorem 5.1]) Assume $M$ splits two hyperbolic planes. Let $x$ be a generic point on the component $\partial_{J}(\overline{{\rm Sh}}(M))$ for some isotropic plane $J\subseteq M$ such that $J^{\perp}/J\in\mathbf{Gen}_{0,n-2}(G_{M})$. The image of a finite linear combinations of Heegner divisors $\sum\limits_{\gamma\in G_{M}}\sum\limits_{m\in\mathbb{Q}}a_{m,\gamma}\mathrm{H}_{m,\gamma}$ is trivial in ${\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M),g\cdot x)$ for all $g\in\mathrm{O}(M)$ if and only if
$\sum\limits_{\gamma\in G_{M}}\sum\limits_{m\in\mathbb{Q}}a_{m,\gamma}c_{-m,\gamma}=0$
for all theta series $\sigma^{\ast}\Theta_{J^{\perp}/J,H}=\sum\limits_{\gamma\in G_{M}}\sum\limits_{m\in\mathbb{Q}}c_{m,\gamma}q^{m}\mathfrak{e}_{\gamma}$, where $H$ runs over all harmonic polynomials of degree $2$ and $\sigma$ runs over $\mathrm{Iso}(G_{M},G_{J^{\perp}/J})$.
For simplicity, we say that an even lattice $M$ is of K3 type if $M$ splits two hyperbolic planes and $M\otimes\mathbb{Z}_{p}$ can split three hyperbolic planes for all primes $p$. Then we have
@@
Let $M$ be an even lattice of signature $(2,n)$ of K3 type. Then if $n>10$,
$\dim_{\mathbb{Q}}{\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M))=1.$
@@
As the boundary of ${\rm Sh}(M)$ in $\overline{{\rm Sh}}(M)$ are 1-dimensional and $\dim{\rm Sh}(M)\geq 3$, the pushforward map is an isomorphism
$i_{\ast}:{\rm Cl}_{\mathbb{Q}}({\rm Sh}(M))\cong{\rm Cl}_{\mathbb{Q}}(\overline{{\rm Sh}}(M)).$
By Theorem 5.3, ${\rm Cl}_{\mathbb{Q}}(\overline{{\rm Sh}}(M))$ is spanned by the Zariski closure of linear combinations of Heegner divisors. One can view ${\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M))$ as a subspace of ${\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))$. For any element $D\in{\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M))$, it can be written as a linear combination $\sum\limits_{\gamma\in G_{M}}\sum\limits_{m\in\mathbb{Q}}a_{m,\gamma}\mathrm{H}_{m,\gamma}$. Since $D$ is $\mathbb{Q}$-Cartier, its image in each local Picard group is a torsion element. By Proposition 6.1 and Theorem 6.2, we have
$\sum\limits_{\gamma\in G_{M}}\sum\limits_{m\in\mathbb{Q}}a_{m,\gamma}c_{-m,\gamma}=0$
for all cusp forms $f=\sum\limits_{\gamma\in G_{M}}\sum\limits_{m\in\mathbb{Q}}c_{m,\gamma}q^{m}\mathfrak{e}_{\gamma}\in\mathrm{Cusp}^{\theta}_{\frac{2+n}{2}}(\rho_{M})$. Applying Theorem 4.5 to $J^{\perp}/J$ for an isotropic plane $J\subset M$ such that $J^{\perp}/J\in\mathbf{Gen}_{0,n-2}(G_{M})$, the space $\mathrm{Cusp}^{\theta}_{\frac{2+n}{2}}(\rho_{M})$ above is actually $\mathrm{Cusp}_{\frac{2+n}{2}}(\rho_{M})$. Now Theorem 5.4 implies that $D$ is proportional to the Hodge line bundle. This proves the theorem. ∎
More generally, if $\Gamma\subset\mathrm{O}(M)(\mathbb{Q})$ is an arithmetic subgroup containing $\widetilde{\mathrm{O}}(M),$denote by ${\rm Sh}_{\Gamma}(M):=\Gamma\backslash\mathbf{D}$ the Shimura variety with respect to $\Gamma$. Let $\overline{{\rm Sh}}_{\Gamma}(M)$ be the Baily-Borel compactification. Then we have
@@
With notations and assumptions as above, if $n>10$, $\dim_{\mathbb{Q}}{\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}_{\Gamma}(M))=1.$
@@
Note that the natural projection
$\overline{{\rm Sh}}(M)\rightarrow\overline{{\rm Sh}}_{\Gamma}(M)$
is a finite surjective morphism. By [3, Corollary 3.8], the first Chern class map induces an isomorphism
${\rm Pic}_{\mathbb{Q}}({\rm Sh}_{(\Gamma)}(M))\cong\mathrm{H}^{2}({\rm Sh}_{(\Gamma)}(M),\mathbb{Q}).$
Thus by the projection formula, the pullback homomorphism ${\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}_{\Gamma}(M))\rightarrow{\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M))$ is injective. Since trivially $\dim_{\mathbb{Q}}{\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}_{\Gamma}(M))\geq 1$, the result now follows from Theorem 6.3. ∎
6.3 Proof of Theorem 1.1
Theorem 6.5 (Theorem 1.1).
${\rm Pic}(\overline{{\mathscr{F}}}_{g})\cong\mathbb{Z}$, which is spanned by an integral multiple of the extended Hodge line bundle $\overline{\lambda}$.
Proof.
This is the case when $M=\Lambda_{g}$. As ${\rm rank}_{p}(G_{\Lambda_{g}})\leq 1$, by Remark 4.6, the conditions in Theorem 6.3 is automatically satisfied. Hence we have $\dim_{\mathbb{Q}}{\rm Pic}_{\mathbb{Q}}(\overline{{\mathscr{F}}}_{g})=1$. As shown in [14], ${\rm Pic}({\mathscr{F}}_{g})$ is torsion-free, then ${\rm Pic}(\overline{{\mathscr{F}}}_{g})$ is also torsion-free as ¯$\overline{{\mathscr{F}}}_{g}\backslash{\mathscr{F}}_{g}$ has codimension $>2$.
It suffices to show that $\mathrm{{\rm Cl}}(\overline{{\mathscr{F}}}_{g})\cap{\rm Pic}_{\mathbb{Q}}(\overline{{\mathscr{F}}}_{g})\subset\mathrm{{\rm Cl}}_{\mathbb{Q}}(\overline{{\mathscr{F}}}_{g})$ is generated by the extended Hodge line bundle $\overline{\lambda}$. To see this, let $\mathcal{S}\to\mathbb{P}^{1}$ be a family of unigonal K3 surfaces of genus $g$ constructed in [22, §4.1.2]:
[…]
$F=\varphi^{-1}(\mathrm{line}),\quad A=\mathbb{P}(\mathcal{O}_{\mathbb{P}^{2}}(4))$
Let $B\in|3A+12F|$ be generic. In particular, it is smooth and it does not intersect $A$. Let $\pi:\widetilde{Y}\to Y$ be the double cover branched over $A+B$. We take a generic member
$\mathcal{S}\in|\mathcal{O}_{\widetilde{Y}}(\pi^{\ast}F)\boxtimes\mathcal{O}_{\mathbb{P}^{1}}(1)|.$
The resulting family $\mathcal{S}\to\mathbb{P}^{1}$ is a family of double cover of the Hirzebruch surface $\Sigma:=\mathbb{P}(\mathcal{O}_{\mathbb{P}^{1}}(4)\oplus\mathcal{O}_{\mathbb{P}^{1}})$ branched over $-2K_{\Sigma}$. The fiber of $\mathcal{S}\to\mathbb{P}^{1}$ is a unigonal K3 surface with at worst A-D-E singularity and admits a polarization of genus $g$, induced by the restriction of $\frac{1}{2}\pi^{\ast}A+g\pi^{\ast}F$.
One can view ${\mathscr{F}}_{g}$ as the coarse moduli space of primitive polarized K3 surfaces of genus $g$ with at worst A-D-E singularities, then we obtain a morphism
$\psi:\mathbb{P}^{1}\to{\mathscr{F}}_{g}\subseteq\overline{{\mathscr{F}}}_{g}$ (6.5)
By using the Grothendieck-Riemann-Roch theorem, Laza and O’Grady have computed that the degree of $\psi^{\ast}(\overline{\lambda}|_{{\mathscr{F}}_{g}})=\psi^{\ast}(\lambda)$ is $1$. This implies that $\overline{\lambda}$ is primitive in $\mathrm{{\rm Cl}}(\overline{{\mathscr{F}}}_{g})\cap{\rm Pic}_{\mathbb{Q}}(\overline{{\mathscr{F}}}_{g})$ and hence
$\mathrm{{\rm Cl}}(\overline{{\mathscr{F}}}_{g})\cap{\rm Pic}_{\mathbb{Q}}(\overline{{\mathscr{F}}}_{g})=\langle\overline{\lambda}\rangle.$
@@
Appendix A A vector-valued Siegel-Weil formula
Following [25] and [34], we introduce the Siegel-Weil formula for metaplectic groups. For simplicity, let us assume $M$ is a negative definite even lattice and $V=M\otimes\mathbb{Q}$. Let ${\mathrm{Sp}}_{2d}$ be the standard symplectic $\mathbb{Z}$-group scheme and ${\rm Mp}_{2d}(\mathbb{A})\stackrel{{\scriptstyle\pi}}{{\rightarrow}}{\mathrm{Sp}}_{2d}(\mathbb{A})$ be the metaplectic double cover. Then the inclusion ${\mathrm{Sp}}_{2d}(\mathbb{Q})\hookrightarrow{\mathrm{Sp}}_{2d}(\mathbb{A})$ can be uniquely extended to an inclusion

### v2

@@ 6.2. Picard Group of $\overline{{\rm Sh}}_{\Gamma}(M)$
Let $M$ be an even lattice of signature $(2,n)$ with $n>8$ and $\Gamma\leq\widetilde{\mathrm{O}}(M)$. Assume that $M$ is K3 type or $p$-elementary for some prime $p$ and splits one hyperbolic plane. Then
$\dim_{\mathbb{Q}}{\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}_{\Gamma}(M))^{\rm Heegner}=1.$
Proof.
Since we assume that $M$ splits a hyperbolic plane, by Proposition 5.13, it suffices to show
${\mathrm{Cusp}}_{k}^{\mathrm{obs}}(\rho_{M}^{*})={\mathrm{Cusp}}_{k}(\rho_{M}^{*}),$
for $k=\frac{2+n}{2}$.
Case 1: $M$ is K3 type.
By Proposition 6.1, for any lattice $L^{\prime}\in\mathbf{Gen}_{0,n-2}(G_{M})$, there exists an isotropic plane $J\subset M$ such that $J^{\perp}/J\cong L^{\prime}$. Since $\rho_{L^{\prime}}\cong\rho_{M}$, applying Theorem 4.5 to $L^{\prime}$ yields
${\mathrm{Cusp}}^{\theta}_{\frac{2+n}{2}}(\rho_{M}^{*})={\mathrm{Cusp}}^{\mathrm{obs}}_{\frac{2+n}{2}}(\rho_{M}^{*})={\mathrm{Cusp}}_{\frac{2+n}{2}}(\rho_{M}^{*}).$
Case 2: $M$ is $2$-elementary.
Let $G_{M}\cong(\mathbb{Z}/2\mathbb{Z})^{a}$ with $a\leq n$. Due to the classification of $2$-elementary lattices, $M$ is of K3 type when $a\leq n-4$. Thus we assume $a>n-4$, implying $a\in\{n-2,n\}$. In this situation, by [Ni79, Thm. 1.10.1 & 1.14.2], there exists $N\in\{1,2\}$ such that $M$ decomposes as:
$M=U(N)\oplus U(2)\oplus L$
where $L$ is negative definite of rank $n-2$ with $G_{L}\cong(\mathbb{Z}/2\mathbb{Z})^{n-4}$. By [Ni79, Cor. 1.9.3], $L$ splits a hyperbolic plane locally, satisfying the assumptions of Theorem 4.5, hence:
${\mathrm{Cusp}}_{k}^{\theta}(\rho_{L}^{*})={\mathrm{Cusp}}_{k}(\rho_{L}^{*}).$
For any $L^{\prime}\in\mathbf{Gen}_{0,n-2}(G_{L})$, Proposition 6.1 yields an admissible isotropic plane $J\subseteq M$ with $J^{\perp}/J\cong L^{\prime}$. Moreover, there exists a natural correspondence between:
  • Admissible isotropic planes $J\subset M$
  • Non-characteristic isotropic subgroups $H\leq G_{M}$ (i.e., containing no nonzero characteristic elements) satisfying $H^{\perp}/H\cong G_{L}$
This consideration for characteristic elements is necessary due to the failure of Witt’s extension theorem in characteristic $2$. Specifically, each admissible $J$ determines a non-characteristic isotropic subgroup $H_{J}\leq G_{M}$ with $H_{J}^{\perp}/H_{J}\cong G_{L}$. Conversely, let $H_{J_{0}}$ denote the isotropic subgroup associated to the standard isotropic plane
$J_{0}\subseteq U(N)\oplus U(2).$
By [Ni79, Thm. 3.6.3 & Lem 3.9.1], any non-characteristic isotropic subgroup $H\leq G_{M}$ with $H^{\perp}/H\cong G_{L}$ is conjugate to $H_{J_{0}}$ under some $\mathrm{O}(M)$-transformation (See also Remark 6.5). This proves $H$ is associated with an admissible isotropic plane.
This correspondence implies:
$\uparrow_{H}\bigl{(}{\mathrm{Cusp}}_{k}(\rho_{H^{\perp}/H}^{*})\bigr{)}=\uparrow_{H}\bigl{(}{\mathrm{Cusp}}^{\theta}_{k}(\rho_{H^{\perp}/H}^{*})\bigr{)}\subseteq{\mathrm{Cusp}}^{\mathrm{obs}}_{k}(\rho_{M}^{*})$
for all such non-characteristic isotropic subgroups $H\leq G_{M}$ and it remains to show that the spaces $\uparrow_{H}\bigl{(}{\mathrm{Cusp}}_{k}(\rho_{H^{\perp}/H}^{*})\bigr{)}$ span ${\mathrm{Cusp}}_{k}(\rho_{M}^{*})$.
By [Ma24a, Thm. 5.1], any $f\in{\mathrm{Cusp}}_{k}(\rho_{M}^{*})$ can be written as
$f=\sum_{H}\uparrow_{H}(f_{H})$ (6.2)
where the sum ranges over all isotropic subgroups $H\leq G_{M}$ such that $H^{\bot}/H\cong(\mathbb{Z}/2\mathbb{Z})^{b}$ with
$b\leq\begin{cases}6&\text{if $n$ is even}\\ 5&\text{if $n$ is odd}\end{cases}$
and $H^{\bot}/H$ has coparity $1$ if $b=6$ with suitable $f_{H}\in{\mathrm{Cusp}}_{k}(\rho_{H^{\bot}/H}^{*})$.
Subcase (i): $a=n-2$. Since $a=n-2>6$, for any $H$ appearing in (6.2) we have $|H|>1$. If $|H|>2$, then $H$ contains a subgroup $\tilde{H}\leq H$ of order $2$ which is non-characteristic. Now assume $|H|=2$. When $a$ is odd, $H$ must be non-characteristic. When $a$ is even, since $a\geq 8$, it follows that $|H^{\bot}/H|=2^{6}$ and therefore, $H^{\bot}/H$ has coparity $1$. Hence, $H$ is non-characteristic. In summary, we find that any $H$ appearing in (6.2) contains a subgroup $\tilde{H}\leq H$ with $|\tilde{H}|=2$ which is non-characteristic. Thus, by the transitivity of the isotropic lifts
$f=\sum_{H}\uparrow_{\tilde{H}}(f_{\tilde{H}})$
where $f_{\tilde{H}}=\uparrow_{H/\tilde{H}}(f_{H})\in{\mathrm{Cusp}}_{k}(\rho_{\tilde{H}^{\bot}/\tilde{H}}^{*})$, which proves the claim.
Subcase (ii): $a=n$. Similarly to before, since $a=n>8$, for any $H$ appearing in (6.2) we have $|H|>2$. If $|H|>4$, then $H$ contains a subgroup $\tilde{H}\leq H$ of order $4$ which is non-characteristic. If $|H|=4$, it must be non-characteristic by an analogous argument as before. Let for any $H$ appearing in (6.2) $\tilde{H}\leq H$ be a subgroup with $|\tilde{H}|=4$ which is non-characteristic. Then,
$f=\sum_{H}\uparrow_{\tilde{H}}(f_{\tilde{H}})$
where $f_{\tilde{H}}=\uparrow_{H/\tilde{H}}(f_{H})\in{\mathrm{Cusp}}_{k}(\rho_{\tilde{H}^{\bot}/\tilde{H}}^{*})$.
Case 3: $M$ is $p$-elementary for $p$ an odd prime.
The proof follows a similar strategy to Case 2. The advantage is that Witt’s theorem holds, while the difficulty lies in the more intricate classification of $p$-elementary lattices.
Note that $n$ must be even and $n\geq 10$. Let $G_{M}\cong(\mathbb{Z}/p\mathbb{Z})^{a}$ with $a\leq n$. By [Ni79, Cor. 1.9.3 and 1.13.5], $M$ is of K3 type if $a\leq n-5$. Thus, we assume $a\geq n-4$. Again, we claim that there exists $N\in\{1,p\}$ such that $M$ decomposes as:
$M=U(N)\oplus U(p)\oplus L$
where $L$ is negative definite of rank $n-2$ and splits locally a hyperbolic plane. This is essentially [Ni79, Cor. 1.10.2 & Theorem 1.14.2]. Specifically,
  • If $a\leq n-3$, then $N=1$ and $L$ splits locally a hyperbolic plane by [Ni79, Cor. 1.9.3].
  • If $a>n-3$, then $N=p$. Since $M$ splits globally a hyperbolic plane and $\ell(G_{U(N)\oplus U(p)})={\rm rank}(U(N)\oplus U(p))$, $L$ splits locally a hyperbolic plane.
Here we shall remark that a negative definite $p$-elementary lattice $L$ does not necessarily locally split a hyperbolic plane when $\ell(G_{L})\leq\mathrm{rank}~{}L-2$.
Theorem 4.5 then implies:
${\mathrm{Cusp}}_{k}^{\theta}(\rho_{L}^{*})={\mathrm{Cusp}}_{k}(\rho_{L}^{*}).$
As in Case 2, one can use Witt’s extension theorem to obtain a correspondence between admissible isotropic planes $J\subset M$ and isotropic subgroups $H\leq G_{M}$ satisfying $H^{\perp}/H\cong G_{L}$. Hence, it suffices to show that the spaces
$\uparrow_{H}\bigl{(}{\mathrm{Cusp}}_{k}(\rho_{H^{\perp}/H}^{*})\bigr{)}\subseteq{\mathrm{Cusp}}^{\mathrm{obs}}_{k}(\rho_{M}^{*}),$
where $H\leq G_{M}$ ranges over isotropic subgroups with $H\cong\mathbb{Z}/N\mathbb{Z}\oplus\mathbb{Z}/p\mathbb{Z}$, span ${\mathrm{Cusp}}_{k}(\rho_{M}^{*})$.
As before, any $f\in{\mathrm{Cusp}}_{k}(\rho_{M}^{*})$ decomposes as:
$f=\sum_{H}\uparrow_{H}(f_{H})$ (6.3)
where the sum is over isotropic subgroups $H\leq G_{M}$ satisfying $H^{\perp}/H\cong(\mathbb{Z}/p\mathbb{Z})^{b}$ with
$b\leq\begin{cases}4&\text{if $a$ is even}\\ 5&\text{if $a$ is odd}\end{cases}$
and $f_{H}\in{\mathrm{Cusp}}_{k}(\rho_{H^{\bot}/H}^{*})$.
Subcase (i): $a\leq n-3$. Since $a\geq n-4\geq 6$, it follows that for any $H$ appearing in (6.3) we have $|H|>1$. Therefore, it contains a subgroup $\tilde{H}\leq H$ with $|\tilde{H}|=p$. Then
$f=\sum_{H}\uparrow_{\tilde{H}}(f_{\tilde{H}})$
where $f_{\tilde{H}}=\uparrow_{H/\tilde{H}}(f_{H})\in{\mathrm{Cusp}}_{k}(\rho_{\tilde{H}^{\bot}/\tilde{H}}^{*})$.
Subcase (ii): $a>n-3$. Since $a>n-3\geq 7$, it follows that for any $H$ appearing in (6.3) we have $|H|>p$. Therefore, it contains a subgroup $\tilde{H}\leq H$ with $|\tilde{H}|=p^{2}$. Then
$f=\sum_{H}\uparrow_{\tilde{H}}(f_{\tilde{H}})$
where $f_{\tilde{H}}=\uparrow_{H/\tilde{H}}(f_{H})\in{\mathrm{Cusp}}_{k}(\rho_{\tilde{H}^{\bot}/\tilde{H}}^{*})$.
Remark 6.5.
For $2$-elementary lattices $M$, Nikulin’s results [Ni79, Thm. 3.6.3 & Lemma 3.9.1] establish that two non-characteristic elements $\gamma,\gamma^{\prime}\in G_{M}$ are $\mathrm{O}(M)$-conjugate if and only if $\mathfrak{q}_{M}(\gamma)=\mathfrak{q}_{M}(\gamma^{\prime})$. Consequently:
  • Any two non-characteristic isotropic subgroups of order $2$ are $\mathrm{O}(M)$-conjugate.
Now consider $M=U(2)\oplus U(2)\oplus L$ with a non-characteristic isotropic subgroup $H\leq G_{M}$ of order $4$. Let $\gamma_{1},\gamma_{2}$ generate $H$ and $e_{i},f_{i}$ the standard basis of $U(2)\oplus U(2)$. By Nikulin’s result:
  • We may assume $\gamma_{1}=e_{1}^{*}$ after $\mathrm{O}(M)$-action;
  • $\gamma_{2}\in G_{U(2)\oplus L}$;
Applying $\mathrm{O}(U(2)\oplus L)$ then maps $\gamma_{2}$ to the generator of the second $U(2)$ factor.
Remark 6.6.
For a $p$-elementary lattice $M$ of signature $(2,n)$, there is a numerical criterion for $M$ to split a hyperbolic plane:
  • If $p=2$, then $M$ splits a hyperbolic plane if and only if $\ell(G_{M})\leq n$.
  • If $p>2$, then $M$ splits a hyperbolic plane if and only if $\ell(G_{M})<n$ or $\ell(G_{M})=n$ and $n\equiv 2\pmod{8}$.
This is due to the work of Nikulin [Ni79].
Corollary 6.7.
Let $M$ be an even lattice of signature $(2,n)$ with $n>8$. We have
$\dim_{\mathbb{Q}}{\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}_{\Gamma}(M))=1$
if either $M$ is of K3 type and $\widetilde{\mathrm{O}}(M)\leq\Gamma$ or $M$ is $p$-elementary and $\Gamma=\Gamma_{0}$.
Proof.
As $\dim M\geq 3$, the natural projection ${\rm Sh}(M)\to{\rm Sh}_{\Gamma}(M)$ preserves the Hodge line bundle and extends to a finite morphism
$\overline{\pi}:\overline{{\rm Sh}}(M)\rightarrow\overline{{\rm Sh}}_{\Gamma}(M),$
between their Baily–Borel compactifications. This induces a commutative diagram
${{\rm Pic}_{\mathbb{Q}}\left(\overline{{\rm Sh}}_{\Gamma}(M)\right)}$${{\rm Pic}_{\mathbb{Q}}\left(\overline{{\rm Sh}}(M)\right)}$${{\rm Pic}_{\mathbb{Q}}\left({\rm Sh}_{\Gamma}(M)\right)}$${{\rm Pic}_{\mathbb{Q}}\left({\rm Sh}(M)\right)}$$\scriptstyle{\overline{\pi}^{*}}$$\scriptstyle{\pi^{\ast}}$
According to Corollary 5.6, in either case, this map will factor through the Heegner divisor subspaces ${\rm Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}(M))^{\rm Heegner}$. Thus we can conclude the assertion by Theorem 6.4. ∎
Proof of Corollary 1.5
Under our hypothesis, the classification result in [AE22] shows that $\Lambda_{\rho}$ is a $2$-elementary lattice and splits a hyperbolic plane. It follows from Corollary 6.7 and Remark 5.1 that ${\rm Pic}((\Gamma_{\rho}\backslash\mathbf{D}_{\rho})^{\rm BB})$ has Picard number one.
6.3. Torsion freeness of the Picard group
We first generalize the results of [DFV24] to Shimura varieties associated with K3-type lattices of large rank.
Definition 6.8.
Let $[{\rm Sh}(M)]=[\widetilde{\mathrm{O}}(M)\backslash\mathbf{D}]$ be the (analytic) quotient stack. It is well-known that $[{\rm Sh}(M)]$ is a smooth Deligne–Mumford stack and its coarse moduli space is ${\rm Sh}(M)$.
We can define an analytic line bundle on $[{\rm Sh}(M)]$ as follows: take the trivial line bundle
$\mathbf{D}\times\mathbb{A}^{1}$
over $\mathbf{D}$ and let $\Gamma=\widetilde{\mathrm{O}}(M)$ act diagonally, where the action on $\mathbb{A}^{1}$ is given by
$A\cdot\lambda\coloneqq\det(A)\lambda.$
The resulting quotient $\mathcal{L}_{M}\coloneqq[\Gamma\backslash\mathbf{D}\times\mathbb{A}^{1}]$ is a line bundle over $[{\rm Sh}(M)]$.
Note that $\mathcal{L}_{M}$ is not trivial because the determinant of an element in $\Gamma$ is not trivial in general.
For any prime $p$, denote by $r_{p}(M)$ the maximal rank of the sublattices $L$ in $M$ such that $\det(L)$ is coprime to $p$. Then we have the following general result.
@@ 6.3. Torsion freeness of the Picard group
Let $M$ be an even lattice of signature $(2,n)$ which splits two hyperbolic lattices. Assume that $r_{2}(M)\geq 6$ and $r_{3}(M)\geq 5$. Then ${\rm Pic}({\rm Sh}(M))$ is torsion free and the torsion part of ${\rm Pic}([{\rm Sh}(M)])$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$, which is generated by $\mathcal{L}_{M}$.
@@
By [GHS09, Thm. 1.7], the abelianisation of $\widetilde{\mathrm{O}}(M)$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$. Following the argument in the proof of [DFV24, Proposition 2.4], the torsion part of $\mathrm{Pic}([\mathrm{Sh}(M)])$ is given by
$\mathrm{Pic}([\mathrm{Sh}(M)])_{\mathrm{torsion}}=\mathbb{Z}\langle\mathcal{L}_{M}\rangle\cong\mathbb{Z}/2\mathbb{Z}.$
This completes the proof of the last assertion.
To show that $\mathrm{Pic}(\mathrm{Sh}(M))$ is torsion-free, we note that by [Ols12, Proposition 6.1], there is an injection
$\mathrm{Pic}(\mathrm{Sh}(M))\hookrightarrow\mathrm{Pic}([\mathrm{Sh}(M)]).$
via the pullback of $[{\rm Sh}(M)]\to{\rm Sh}(M)$. Thus, it suffices to prove that $\mathcal{L}_{M}$ does not descend to $\mathrm{Sh}(M)$. Set $M_{0}=U^{\oplus 2}$ and write $M=M_{0}\oplus N$. Then the natural embedding
$M_{0}\hookrightarrow M$
induces a Cartesian diagram
${\left[\mathrm{Sh}(M_{0})\right]}$${\left[\mathrm{Sh}(M)\right]}$${\mathrm{Sh}(M_{0})}$${\mathrm{Sh}(M).}$
The pullback of $\mathcal{L}_{M}$ to $[{\rm Sh}(M_{0})]$ is $\mathcal{L}_{M_{0}}$. If $\mathcal{L}_{M}$ descends to ${\rm Sh}(M)$, then $\mathcal{L}_{M_{0}}$ also descends to ${\rm Sh}(M_{0})$. However, this contradicts the fact ${\rm Sh}(M_{0})\cong\mathbb{A}^{2}$ and only the trivial line bundle descends to ${\rm Sh}(M_{0})$.
As an immediate consequence, we obtain
@@
Let $M$ be a lattice of K3 type and ${\rm rank}(M)>10$. Then ${\rm Pic}(\overline{{\rm Sh}}(M))\cong\mathbb{Z}$ is spanned by some multiple of $\overline{\lambda}_{M}$.
@@
Since $M$ is of K3 type, it satisfies the conditions $r_{2}(M)\geq 6$ and $r_{3}(M)\geq 5$. By Theorem 6.9, ${\rm Pic}({\rm Sh}(M))$ is torsion-free. As $\overline{{\rm Sh}}(M)\setminus{\rm Sh}(M)$ has codimension $>2$, ${\rm Pic}(\overline{{\rm Sh}}(M))$ is also torsion-free. This implies ${\rm Pic}(\overline{{\rm Sh}}(M))\cong\mathbb{Z}$. ∎
@@
Appendix A A vector-valued Siegel–Weil formula
Following [Ma24] and [MpSW], we introduce the Siegel–Weil formula for metaplectic groups. For simplicity, let us assume $M$ is a positive definite even lattice and $V=M\otimes\mathbb{Q}$. Let ${\mathrm{Sp}}_{2d}$ be the standard symplectic $\mathbb{Z}$-group scheme and ${\rm Mp}_{2d}(\mathbb{A})\stackrel{{\scriptstyle\pi}}{{\rightarrow}}{\mathrm{Sp}}_{2d}(\mathbb{A})$ be the metaplectic double cover. Then the inclusion ${\mathrm{Sp}}_{2d}(\mathbb{Q})\hookrightarrow{\mathrm{Sp}}_{2d}(\mathbb{A})$ can be uniquely extended to an inclusion